How many years has passed if only 3.125% of the original amount of U-238 remains?


1
Expert's answer
2021-04-18T19:32:57-0400

Given

"R=0.03125R_0"

"T_{1\/2}=4.5\u00d710^9 years"

The amount of radioactivity present after time t is

"R=R_0 e^{-\frac{tln2}{t_{1/2}}}"

"0.03125R_0=R_0 e^{-\frac{tln2}{4.5×10^9 years}}"

"t=2.25 \times 10^{10} years"